Tags:
create new tag
view all tags

Cavity BPM Shift

present: Stewart, Jochem, Young-Im

ATF logbook: http://atf.kek.jp/twiki/bin/view/ATFlogbook/Log20130416d

10:00 shift started

  • tunnel access to remove attenuation from QD10BFF
  • Need to tune and calibrate MFB2FF (+- 50 um) , QD10BFF (+-250 um), IPX1, IPX2 (+- 250 um)
  • DR frequency change -4kHz to 714.019MHz
  • intensity about 0.4e10
Operation BPM File name Note
tune MFB2FFx bpmLog_20130416_110403.dat.gz  
tune MFB2FFy bpmLog_20130416_110752.dat.gz Tau fit fail
tune QD10BFFx bpmLog_20130416_111815.dat.gz  
tune QD10BFFy bpmLog_20130416_111939.dat.gz  
tune REFIPX bpmLog_20130416_112157.dat.gz  
svd cal MFB2FFx bpmAllMoverCalibration_20130416_112547.dat.gz looked strange (check) +- 50
svd cal MFB2FFy bpmAllMoverCalibration_20130416_114206.dat.gz 56.68, 2.6908
svd cal QD10BFFx bpmAllMoverCalibration_20130416_114930.dat.gz  
svd cal QD10BFFy bpmAllMoverCalibration_20130416_115309.dat.gz  
fs bump cal IPAx bpmFsBumpCalibration_20130416_115807.dat.gz  
fs bump cal IPBx bpmFsBumpCalibration_20130416_120133.dat.gz  

  • Constant upload off, Sextupoles off( previous SF6FF : 15.855, SD4FF : 28.152, SF5FF : 1.067, SF1FF 3.769, SD0FF 6.361)
  • 500 pulse file : bpmAllLog_20130416_120535.dat.gz
QF19Xx -> about 15% change in pos scale (orig: 1053, new: 1206, repeat: 1160). IQ -rot (orig: 0.44, new: 0.45, repeat 0.45)

QM15FFy -> about 10% change in pos scale

QM14FFx bpmMoverCalibration_20130416_132705.dat.gz -> mover trim problem, redo.

13:30 beam off for magnet power supply switch.

14:40 beam on.

14:55 restart calibrations, start with QM14FFx

intensity 0.2e10, Okugi-san and Naito-san tuning to 0.5e10 continuing calibration. Tuning was done during calibration of QM13FFy to QD10BFFx

QM11FFy: max power interlock -> only 4 svd positions. redone.

We found out that QD2AFFy was not calibrated correctly!--> need to check original calibration

original sextupole positions:

sextupole magnet x y
SF6FF 280.7 -115.0
SF5FF 311.0 9.0
SD4FF 97.6 -133.4
SF1FF 307.7 -146.5
SD0FF 45.8 -191.0

QF7FFy svd crash. redo. (no trace due to 'less' screwing up history) files:

20130416_164956 bpmSvdMoverCal QF7FF y
20130416_165005 bpmLogAlldata bpmAllLog_20130416_164957.dat.gz 20.0
20130416_165332 bpmSvdMoverCal bpmAllMoverCalibration_20130416_165006.dat.gz QF7FF y 20.0 -250.0 250.0 5.0
20130416_165332 bpmMoverCal bpmMoverCalibration_20130416_165006.dat.gz QF7FF y 20.0 -250.0 250.0 5.0

QD2AFFy calibrated and uploaded 1062.18 1.8595

SBand calibrate with +- 250um. small signals

SF1FF x bpmMoverCalibration_20130416_183616.dat.gz
SF1FF y bpmMoverCalibration_20130416_183827.dat.gz
SD0FF x bpmMoverCalibration_20130416_184026.dat.gz
SD0FF y bpmMoverCalibration_20130416_184209.dat.gz

remove 6dB attenuation. recalibrate and upload constants, and change range

BPM log scale IQrot range
SF1FF x bpmMoverCalibration_20130416_184443.dat.gz 2243.39 2.8415 +-250
SF1FF x bpmMoverCalibration_20130416_184636.dat.gz 1978.22   +-500
SF1FF y bpmMoverCalibration_20130416_184824.dat.gz 1368.32 -1.0098 +-500
SD0FF x bpmMoverCalibration_20130416_185021.dat.gz 1404.70 0.1290 +-500
SD0FF y bpmMoverCalibration_20130416_185204.dat.gz 887.13 -0.9512 +-500
FB2FFy bpmMoverCalibration_20130416_185443.dat.gz 131.77 2.6852 +-50
FB2FFx bpmMoverCalibration_20130416_185721.dat.gz 113.45 -0.4043 +-250

save calibration data: * bpmCasr:save bpmCasr_20130416_190033.dat *

PREIP P-IP IPA IPB bump recalibration

IPB-x calibration position scale different from database (3000 vs 6000) redo, same as before. was fine all along? to be checked

IPB-x 1 bpmFsBumpCalibration_20130416_192004.dat.gz

IPB-x 2 bpmFsBumpCalibration_20130416_192526.dat.gz

flatten orbit at IP region with QD0 and QF7FF and take orbit data.

IP region steered resolution data (with QD0 and QF7) looks good!

file note
bpmAllLog_20130416_193622.dat  
bpmAllLog_20130416_193954.dat  
bpmAllLog_20130416_194438.dat increase charge a bit (0.6e10 EXT, 0.3e10 IP)
bpmAllLog_20130416_194821.dat decrease charge a bit (0.4e10 EXT, 0.2e10 IP), but gotten up by tuning of Fukuda-san
bpmAllLog_20130416_195447.dat increase charge again (0.5e10 EXT, 0.3e10 IP)
bpmAllLog_20130416_200134.dat steer orbit around with ZV6X scan
bpmAllLog_20130416_201520.dat steer orbit around with ZH1X scan

*IPA-y and IPB-y redo Calibration, constant upload off

bpmFsBumpCalibration_20130416_211039.dat.gz

bpmFsBumpCalibration_20130416_211345.dat.gz

  • /atf/data/cbpm/raw/bpmAllLog_20130417_180123.dat.gz

Edit | Attach | Watch | Print version | History: r12 < r11 < r10 < r9 < r8 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r12 - 17 Apr 2013 - StewartBoogert

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2022 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback